# Disentangling Knowledge Representations for Large Language Model Editing

### v1

@@ 4.3 Experiment Results
How does DiKE Perform in Preserving Fine-grained knowledge? To avoid performance inflation due to data leakage and to verify the generalization capability of our KRD module, we ensure minimal subject overlap between the KRD training data and the evaluation datasets. Specifically, the subject overlap rates between the KRD module’s training set and the edited samples in FINE-KED and COUNTERFACT are only 1.39% and 6.33%, respectively.
@@
(i) Performance on FINE-KED. Table 2 presents the performance of all editors on FINE-KED. From the results, we can draw the following observations: (1) DiKE excels in all evaluations of Relational Locality while maintaining high editing success rate. Specifically, DiKE performs enhancements up to 8.3% on the Relational Locality metric over the baseline model. This demonstrates that fine-grained disentanglement and constraint of target knowledge can effectively mitigate negative impacts on fine-grained unrelated knowledge without compromising editing performance. (2) ROME, MEMIT and AlphaEdit generally perform worse on the Relational Locality compared to DiKE. MEND and FT, demonstrate better performance than ROME and MEMIT on the Easy and Middle levels of FINE-KED. However, their performance declines significantly on the Hard level of the dataset. We believe this is because these models directly manipulate the subject or relation representation to perform parameter updates, which can easily have negative impacts on fine-grained unrelated knowledge that is entangled within the parameters and representations. Although AlphaEdit introduces null-space constraints to preserve unrelated knowledge, the constraint is applied at a coarse-grained level and thus struggles to prevent interference with fine-grained irrelevant knowledge.
@@
(3) ROME-C and MEMIT-C, which incorporate additional relational constraints, improve upon their original versions but still fall short of DiKE. This suggests that merely adding constraint samples is insufficient to reliably preserve fine-grained unrelated knowledge. Furthermore, the requirement to manually construct extra constraints for each editing instance introduces substantial overhead. In contrast, DiKE, with its KRD moudle, requires only one-time training process and can be directly applied during subsequent edits without retraining, significantly enhancing editing efficiency.
(ii) Performance on COUNTERFACT. Table 3 presents the performance of all editors on COUNTERFACT. The results show that DiKE achieves competitive results across other key editing evaluation metrics, such as Paraphrase Score and Neighborhood Score. These findings suggest that the disentanglement-based approach, DiKE, not only effectively preserves fine-grained irrelevant knowledge but also maintains strong generalization ability and better retention of unrelated neighborhood knowledge.
@@
We observe that DiKE achieves an Efficacy Score close to 100% across most models and batch settings, while maintaining the highest Relational Locality Score. ROME exhibits a significant decline in performance as the batch size increases. Although MEMIT and AlphaEdit introduce mechanisms for batch editing and unrelated knowledge preservation, respectively, their improvements in Relational Locality under larger batch settings remain significantly lower than those of DiKE. A possible reason for this advantage is that DiKE explicitly disentangles the subject representation with respect to different relations, thereby isolating relation-specific knowledge within the subject. As a result, edits to one relation incur minimal interference with other knowledge associated with the same subject. In contrast, methods like ROME, MEMIT, and AlphaEdit directly update the subject representation, where relationally entangled knowledge may conflict during editing—leading to less precise updates and reduced effectiveness. Furthermore, AlphaEdit only enforces constraints on coarse-grained unrelated knowledge, making it less effective at preserving fine-grained relational distinctions within the subject representation.
@@ 5 Related Work
Knowledge editing aims to inject or modify knowledge in LLMs while minimizing interference with unrelated information. We focus on parameter-modifying methods, which fall into three categories: fine-tuning-based approaches (e.g., RECT Gu et al. (2024)) that update relevant parameters with constraints; meta-learning methods (e.g., MEND Mitchell et al. (2022), MALMEN Tan et al. (2024)) that use hypernetworks to generate edits; and locate-then-edit frameworks (e.g., ROME Meng et al. (2022), MEMIT Meng et al. (2023), AlphaEdit Fang et al. (2025)) that identify and modify knowledge-bearing parameters. More detailed related work is provided in Appendix C.
@@ 6 Conclusion
Appendix A Limitations
@@
Second, the Knowledge Representation Disentanglement (KRD) module in DiKE requires additional data for pretraining. Although the trained KRD module demonstrates strong generalization and can be reused across editing instances without retraining, the quality of the disentangled representations may still be sensitive to the quality and diversity of the pretraining data, as well as the selection of hyperparameters.
@@
Third, the scope of our evaluation benchmark, FINE-KED, is currently limited to measuring editing success and the preservation of fine-grained irrelevant knowledge. While it provides a targeted assessment of editing success and preservation of fine-grained irrelevant knowledge, these dimensions alone are insufficient to fully characterize the broader capabilities required in knowledge editing. In particular, many real-world editing scenarios involve additional challenges such as generalization to paraphrases, multi-hop reasoning, and robustness to distribution shifts. Therefore, FINE-KED should be viewed as a complementary benchmark rather than a standalone evaluation. In future work, we plan to further expand FINE-KED to better capture the diverse challenges in model editing and provide a more complete assessment of editing performance.
Appendix B Impact Statement
This paper presents work whose goal is to advance the field of knowledge editing for LLMs. Our method, DiKE, improves the precision of knowledge updates while minimizing unintended interference with fine-grained unrelated knowledge. By enhancing the reliability and controllability of LLMs, our approach contributes to more trustworthy and adaptive AI systems.
While knowledge editing has broad applications, including AI alignment, fact-checking, and continual learning, we acknowledge potential ethical considerations, such as the responsible modification of AI-generated knowledge. We encourage future work to explore safeguards against misuse and ensure that knowledge editing techniques are deployed in alignment with ethical AI principles.
Appendix C Related Work
In this section, we introduce the related work on knowledge editing, which aims to inject new knowledge into LLMs or modify their internal knowledge, while minimizing unintended changes to unrelated knowledge. This study focuses on parameter-modifying methods, which can be broadly categorized into three groups: Fine-tuning-based methods [Gu et al., 2024, Yu et al., 2024, Ni et al., 2024] utilize efficient parameter-tuning techniques to update model knowledge. To alleviate issues such as overfitting, these methods typically introduce additional constraints to preserve unrelated knowledge. For example, RECT [Gu et al., 2024] injects new knowledge by selecting and fine-tuning the top-$k$ parameters most relevant to the target, while simultaneously constraining the magnitude of updates to reduce interference with other knowledge. Meta-learning-based methods employ a hypernetwork to generate editing-specific parameter updates. MEND [Mitchell et al., 2022] uses a low-rank gradient decomposition and a lightweight hypernetwork to transform fine-tuning gradients into weight updates. MALMEN [Tan et al., 2024] extends MEND to batch editing by formulating the update process as a least-squares optimization problem. Locate-then-edit Methods perform editing by first identifying model parameters associated with the target knowledge and then applying targeted updates [Meng et al., 2023, Gupta et al., 2024, Zhang et al., 2024, Fang et al., 2025, Zhang et al., 2025]. Early work such as Knowledge Neurons [Dai et al., 2022], proposed a knowledge attribution method to identify relevant neurons. However, this approach exhibits limitations in precisely adjusting model weights. Subsequently, ROME [Meng et al., 2022] treats the weights of the FFN layers as a form of linear associative memory and updates specific layers to encode new knowledge. MEMIT [Meng et al., 2023] extends ROME by enabling large-scale knowledge editing through shared updates across multiple layers, thereby mitigating interference with previously edited layers. AlphaEdit [Fang et al., 2025] extends this line of work by introducing null-space constraints to prevent updates from affecting unrelated knowledge. However, its constraints are applied at a coarse-grained level, making it less effective at preserving fine-grained irrelevant knowledge that may be entangled with the target in the representation space.
@@ Appendix C Related Work
Appendix D Derivation Details for Fine-Grained Irrelevant Knowledge Preservation
@@
We provide a detailed derivation of Equation (17) from Section 3.2, which enforces the consistency of the disentangled target-knowledge-unrelated representations before and after editing:
@@
$\hat{\mathbf{W}}=\underset{\hat{\mathbf{W}}}{\operatorname{argmin}}\bigg{(}$ $\underbrace{\left\|\hat{\mathbf{W}}\mathbf{k}_{*}-\mathbf{v}_{*}\right\|_{F}^{2}}_{\textbf{Target knowledge editing}}+\underbrace{\left\|\hat{\mathbf{W}}\mathbf{K}_{0}-\mathbf{V}_{0}\right\|_{F}^{2}}_{\textbf{Coarse-grained irrelevant knowledge preserving}}$
@@
Since $\mathbf{W}$ is assumed to be the optimal least-squares solution for memorizing a mapping from a previous set of keys $\mathbf{K}_{0}$ to values $\mathbf{V}_{0}$, it satisfies the normal equation [Meng et al., 2022]:
@@
Appendix E Details of Datasets & Evaluation Metrics
E.1 Details of FINE-KED
@@
Following RippleEdits [Cohen et al., 2024], we first sample a set of subjects and their corresponding knowledge triples $T_{s}=\{(s,r_{i},o_{i})|i=1,2,...\}$. Specifically, we select entities as subjects if their corresponding Wikipedia pages ranked within the top-$1000$ most viewed pages for at least one month during 2020 to 2022. For each relation $r$, we collect a set of objects $O_{r}$, comprising all objects from knowledge triples sharing the relation $r$. For every subject, we randomly select one triple $(s,r,o)$ as the edit prompt and sample a target object $o^{*}\neq o$ from $O_{r}$. To construct the neighborhood prompt, we use the GPT-J model to filter out knowledge already correctly recalled by the model (i.e., triples where the model’s accuracy exceeds than 80%) and then randomly select one remaining $(s,r^{\prime},o^{\prime})$ as the fine-grained neighborhood prompt.
To comprehensively evaluate the effectiveness of editing methods in preserving fine-grained irrelevant knowledge, we classify all neighborhood prompts into three difficulty levels based on their relational similarity to the edit prompt: Easy, Middle and Hard. Specifically, we prompt Mistral-7B-Instruct-v0.3 (see Table E.1 for details) to evaluate relational similarities on a scale from 0 (completely unrelated) to 10 (highly related). We define the categories as follows: Easy (0–3), Middle (4–6), and Hard (7–10).
@@
- Easy Level 1501
- Middle Level 827
- Hard Level 757
@@
Evaluation Criteria: 1. Very Similar (score 7-10): If the two relationships describe entities of the same type or are very close, differing only in details (e.g., "capital" and "largest city"). 2. Moderately Similar (score 4-6): If the two relationships describe different types of entities but are in the same domain or background, or have some overlap in the entities they describe (e.g., "mother" and "place of birth"). 3. Completely Not Similar (score 0-3): If the two relationships describe entirely different types of entities from different domains or categories, with almost no relation (e.g., "member of sports team" and "head of state"). Example Inputs and Outputs: 1. Example 1: Relationship 1: r1 = currency (currency used by item) Relationship 2: r2 = capital (seat of government of a country, province, state or other type of administrative territorial entity) Output: Score: 4 Explanation: These two relationships describe different types of entities—one is about currency, and the other is about political/geographic entities. Although both relate to countries, they differ significantly in their descriptions, so they are moderately similar. 2. Example 2: Relationship 1: r1 = place of birth (most specific known birth location of a person, animal or fictional character) Relationship 2: r2 = spouse (the subject has the object as their spouse (husband, wife, partner, etc.)) Output: Score: 2 Explanation: These two relationships describe completely different things—one is about a birthplace, and the other is about a marriage relationship. There is little to no overlap, so they are completely not similar. 3. Example 3: Relationship 1: r1 = head of state (the chief public representative of a country) Relationship 2: r2 = head of government (the person in charge of running the government of a country) Output: Score: 9 Explanation: These two relationships describe very similar entities—both refer to the highest leaders of a country, with "head of state" focusing on ceremonial roles and "head of government" focusing on executive responsibilities. They are very similar. Your task: Relationship 1: r1 = {relation_A} Relationship 2: r2 = {relation_B}
@@
[…]
@@ E.2 Details of COUNTERFACT
To evaluate COUNTERFACT across all editors, we adopt three widely used metrics [Meng et al., 2022, 2023], Efficacy, Paraphrase Score and Neighborhood Score. Each metric is calculated as follows:
@@
Appendix F Baselines
Our experiments are conducted on GPT-2 XL (1.5B) [Radford et al., 2019], GPT-J (6B) [Wang and Komatsuzaki, 2021] and LLaMA3 (8B) [Dubey et al., 2024]. We compare the DiKE against the following state-of-the-art editing techniques:
  • Constrained Fine-Tuning (FT) [Zhu et al., 2020], which directly fine-tunes specific layers of LLM’s parameters using gradient descent while applying regularization constraints to prevent catastrophic forgetting;
  • MEND [Mitchell et al., 2022], a gradient-based low-rank decomposition method that employs a hypernetwork to perform edits;
  • ROME [Meng et al., 2022], which assumes that knowledge in LLMs is stored in FFN modules and performs edits by optimizing and updating specific FFN layers to insert knowledge;
  • MEMIT [Meng et al., 2023], an extension of ROME designed specifically for batch editing tasks by editing a sequence of FFN layers.
  • AlphaEdit [Fang et al., 2025], a null-space projection method designed to better preserve unrelated knowledge during editing by constraining updates orthogonal to preserved information.

### v2

@@ 4.3 Experiment Results
How does DiKE Perform in Preserving Fine-grained knowledge? To avoid performance inflation due to data leakage and to verify the generalization capability of our KRD module, we ensure minimal subject overlap between the KRD training data and the evaluation datasets. Specifically, the subject overlap rates between the KRD module’s training set and the edited samples in FINE-KED and COUNTERFACT are only 1.39% and 6.33%, respectively. A detailed investigation of the generalization behavior of the KRD module is provided in Appendix G.6.
@@
(i) Performance on FINE-KED. Table 2 presents the performance of all editors on FINE-KED. From the results, we can draw the following observations: (1) DiKE excels in all evaluations of Relational Locality while maintaining high editing success rate. Specifically, DiKE performs enhancements up to 8.3% on the Relational Locality metric over the baseline model. This demonstrates that fine-grained disentanglement and constraint of target knowledge can effectively mitigate negative impacts on fine-grained unrelated knowledge without compromising editing performance. (2) ROME, MEMIT and AlphaEdit generally perform worse on the Relational Locality compared to DiKE. MEND and FT, demonstrate better performance than ROME and MEMIT on the Easy and Middle levels of FINE-KED. However, their performance declines significantly on the Hard level of the dataset. We believe this is because these models directly manipulate the subject or relation representation to perform parameter updates, which can easily have negative impacts on fine-grained unrelated knowledge that is entangled within the parameters and representations. Although AlphaEdit introduces null-space constraints to preserve unrelated knowledge, the constraint is applied at a coarse-grained level and thus struggles to prevent interference with fine-grained irrelevant knowledge. (3) ROME-C and MEMIT-C, which incorporate additional relational constraints, improve upon their original versions but still fall short of DiKE. This suggests that merely adding constraint samples is insufficient to reliably preserve fine-grained unrelated knowledge. Furthermore, the requirement to manually construct extra constraints for each editing instance introduces substantial overhead. In contrast, DiKE, with its KRD moudle, requires only one-time training process and can be directly applied during subsequent edits without retraining, significantly enhancing editing efficiency.
(ii) Performance on COUNTERFACT. Table 3 presents the performance of all editors on COUNTERFACT. The results show that DiKE achieves competitive results across other key editing evaluation metrics, such as Paraphrase Score and Neighborhood Score.
@@
These findings suggest that the disentanglement-based approach, DiKE, not only effectively preserves fine-grained irrelevant knowledge but also maintains strong generalization ability and better retention of unrelated neighborhood knowledge.
@@
We observe that DiKE achieves an Efficacy Score close to 100% across most models and batch settings, while maintaining the highest Relational Locality Score. By contrast, ROME exhibits a sharp degradation as batch size increases. Although MEMIT and AlphaEdit incorporate mechanisms for batch editing and irrelevant knowledge preservation, respectively, their improvements in Relational Locality under larger batch settings remain significantly lower than those of DiKE. This advantage arises because DiKE explicitly disentangles subject representations across different relations, thereby isolating relation-specific knowledge and minimizing interference. As a result, edits to one relation have little impact on other knowledge associated with the same subject. By contrast, ROME, MEMIT, and AlphaEdit directly update entangled subject representations, where overlapping relational information leads to conflicts, imprecise updates, and reduced effectiveness. Furthermore, AlphaEdit constrains only coarse-grained irrelevant knowledge, making it less effective at preserving fine-grained relational distinctions within subject representations. We also find that MEND suffers from a significant performance drop in this setting, which may be attributed to its reliance on a hypernetwork to predict parameter updates based on gradients. When editing a batch of knowledge instances sharing the same subject, the predicted gradients are more likely to conflict with each other, resulting in editing failures.
@@ 5 Related Work
In this section, we introduce the related work on knowledge editing, which aims to inject new knowledge into LLMs or modify their internal knowledge, while minimizing unintended changes to unrelated knowledge. This study focuses on parameter-modifying methods, which can be broadly categorized into three groups:
Fine-tuning-based methods (Gu et al., 2024; Yu et al., 2024; Ni et al., 2024) utilize efficient parameter-tuning techniques to update model knowledge. To alleviate issues such as overfitting, these methods typically introduce additional constraints to preserve unrelated knowledge. For example, RECT (Gu et al., 2024) injects new knowledge by selecting and fine-tuning the top-$k$ parameters most relevant to the target, while simultaneously constraining the magnitude of updates to reduce interference with other knowledge.
Meta-learning-based methods employ a hypernetwork to generate editing-specific parameter updates. MEND (Mitchell et al., 2022) uses a low-rank gradient decomposition and a lightweight hypernetwork to transform fine-tuning gradients into weight updates. MALMEN (Tan et al., 2024) extends MEND to batch editing by formulating the update process as a least-squares optimization problem.
Locate-then-edit Methods perform editing by first identifying model parameters associated with the target knowledge and then applying targeted updates (Meng et al., 2023; Gupta et al., 2024; Zhang et al., 2024; Fang et al., 2025; Zhang et al., 2025b; a). Early work such as Knowledge Neurons (Dai et al., 2022), proposed a knowledge attribution method to identify relevant neurons. However, this approach exhibits limitations in precisely adjusting model weights. Subsequently, ROME (Meng et al., 2022) treats the weights of the FFN layers as a form of linear associative memory and updates specific layers to encode new knowledge. MEMIT (Meng et al., 2023) extends ROME by enabling large-scale knowledge editing through shared updates across multiple layers, thereby mitigating interference with previously edited layers. AlphaEdit (Fang et al., 2025) extends this line of work by introducing null-space constraints to prevent updates from affecting unrelated knowledge. However, its constraints are applied at a coarse-grained level, making it less effective at preserving fine-grained irrelevant knowledge that may be entangled with the target in the representation space.
@@ 6 Conclusion
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Nos. 62502286, 62472261, 62576339, 62372275). It was also Funded by Basic Research Program of Jiangsu (BK20250429), the Shandong Provincial Natural Science Foundation (ZR2024QF203), the Technology Innovation Guidance Program of Shandong Province (YDZX2024088), the Provincial Key R&D Program of Shandong Province (2024CXGC010108), and the Open Research Fund of the State Key Laboratory of Multimodal Artificial Intelligence Systems.
Ethics Statement
The goal of this work is to advance the field of knowledge editing in LLMs, aiming to improve the precision and reliability of information they contain. Our proposed method, DiKE, is designed to update factual knowledge while minimizing unintended side effects on related information, which we believe is a step toward more controllable and trustworthy AI systems. Such improvements have positive applications in areas like correcting factual inaccuracies, removing harmful or biased content, and keeping models updated with the latest information without costly retraining.
However, we acknowledge that like any technology that allows for the modification of information, knowledge editing tools could be misused. A malicious actor could potentially leverage such techniques to insert subtle misinformation, propaganda, or harmful biases into an LLM. While our work focuses on improving the technical fidelity of edits, it does not in itself prevent such applications. We believe the broader research community must continue to develop robust detection mechanisms and safeguards in parallel with advancements in editing techniques. We encourage the responsible development and deployment of knowledge editing technologies, with a strong emphasis on ethical considerations and safeguards against misuse.
Reproducibility Statement
To ensure the reproducibility of our results, we provide detailed descriptions of our methodology and experimental setup throughout the paper. All experiments were conducted using publicly available Large Language Models: GPT2-XL (1.5B), GPT-J (6B), and LLaMA3 (8B). We evaluate our method on the widely-used COUNTERFACT and MQUAKE-3K benchmarks. For evaluating the preservation of fine-grained irrelevant knowledge, we introduce a new benchmark named FINE-KED, and we will release the dataset publicly upon publication.
Appendix F.1 provides specific implementation details, including key hyperparameters such as learning rates, batch sizes, optimizer settings, and the specific layers targeted for editing in each model. For baseline comparisons, we utilized the experimental framework provided by Meng et al. (2023) to ensure consistency. To further aid reproducibility, we will make our source code and the newly constructed FINE-KED dataset publicly available.
Appendix A The Use of Large Language Models
During the preparation of this work, we use a LLM for assistance with the following tasks: (1) refining the language and improving the clarity of the manuscript, particularly in the Methodology and Results sections; (2) correcting grammatical errors and ensuring stylistic consistency.
Appendix B Limitations & Future Discussion
@@
Second, the scope of our evaluation benchmark, FINE-KED, is currently limited to measuring editing success and the preservation of fine-grained irrelevant knowledge. While it provides a targeted assessment of editing success and preservation of fine-grained irrelevant knowledge, these dimensions alone are insufficient to fully characterize the broader capabilities required in knowledge editing. In particular, many real-world editing scenarios involve additional challenges such as generalization to paraphrases, multi-hop reasoning, and robustness to distribution shifts. Therefore, FINE-KED should be viewed as a complementary benchmark rather than a standalone evaluation. In future work, we plan to further expand FINE-KED to better capture the diverse challenges in model editing and provide a more complete assessment of editing performance.
Finally, we focus specifically on fine-grained irrelevant knowledge that shares the same subject as the edited fact in this study. This choice is motivated by the observation that existing representative editing methods, such as AlphaEdit and ROME, perform poorly on this type of knowledge (as shown in Fig.1), highlighting it as a particularly challenging and underexplored problem. We believe that addressing this category is both meaningful and necessary. Moreover, in structured knowledge editing scenarios, defining fine-grained irrelevant knowledge via shared subjects but differing relations is both intuitive and formally tractable, which facilitates evaluation and methodological design. In future work, we plan to explore other types of fine-grained irrelevant knowledge beyond subject overlap to further broaden the scope and generality of our framework.
@@
Appendix C Derivation Details for Fine-Grained Irrelevant Knowledge Preservation
@@
We provide a detailed derivation of Equation equation 17 from Section 3.2, which enforces the consistency of the disentangled target-knowledge-unrelated representations before and after editing:
@@
[…]
@@
Since $\mathbf{W}$ is assumed to be the optimal least-squares solution for memorizing a mapping from a previous set of keys $\mathbf{K}_{0}$ to values $\mathbf{V}_{0}$, it satisfies the normal equation (Meng et al., 2022):
@@
Appendix D Details of Datasets & Evaluation Metrics
D.1 Details of FINE-KED
@@
Following RippleEdits (Cohen et al., 2024), we first sample a set of subjects and their corresponding knowledge triples $T_{s}=\{(s,r_{i},o_{i})|i=1,2,$...\}. Specifically, we select entities as subjects if their corresponding Wikipedia pages ranked within the top-$1000$ most viewed pages for at least one month during 2020 to 2022. For each relation $r$, we collect a set of objects $O_{r}$, comprising all objects from knowledge triples sharing the relation $r$. For every subject, we randomly select one triple $(s,r,o)$ as the edit prompt and sample a target object $o^{*}\neq o$ from $O_{r}$. To construct the neighborhood prompt, we use the GPT-J model to filter out knowledge already correctly recalled by the model (i.e., triples where the model’s accuracy exceeds than 80%) and then randomly select one remaining $(s,r^{\prime},o^{\prime})$ as the fine-grained neighborhood prompt.
To comprehensively evaluate the effectiveness of editing methods in preserving fine-grained irrelevant knowledge, we classify all neighborhood prompts into three difficulty levels based on their relational similarity to the edit prompt: Easy, Middle and Hard. Specifically, we prompt Mistral-7B-Instruct-v0.3 (see Table F.1 for details) to evaluate relational similarities on a scale from 0 (completely unrelated) to 10 (highly related). We define the categories as follows: Easy (0–3), Middle (4–6), and Hard (7–10). To ensure the reliability of our difficulty-level categorization, we conducted a human evaluation with expert annotators from the field of LLMs. The results confirmed a high level of agreement between the labels generated by our method and the judgments of human experts.
@@
  - Easy Level 1501
  - Middle Level 827
  - Hard Level 757
@@
Edit Prompt {Selma Kurz} was employed in Vienna $\rightarrow$ London.
@@ D.2 Details of COUNTERFACT
To evaluate COUNTERFACT across all editors, we adopt three widely used metrics (Meng et al., 2022; 2023), Efficacy, Paraphrase Score and Neighborhood Score. Each metric is calculated as follows:
@@
D.3 Details of MQuAKE-3K
MQuAKE-3K (Zhong et al., 2023) is a challenging benchmark for evaluating whether models can perform multi-hop reasoning with newly edited knowledge. Each instance consists of multiple single-hop edits, accompanied by questions that require multi-hop reasoning over the updated facts. This setup places stricter demands on edited LLMs, as they must not only memorize new information but also integrate it across reasoning chains. Table 7 provides an example from MQuAKE-3K dataset. To fully exploit the reasoning capabilities of LLMs, we adopt a zero-shot setting for answer generation. Following Zhong et al. (2023), we report the Efficacy Score to measure the accuracy of the post-edit model on the multi-hop question set $P$ about the edit sample: $\mathbb{E}_{q\in Q}[\mathbb{I}[\mathrm{P}(\text{new answer}|q)>\mathrm{P}(\text{original answer}|q)]]$.
Table 7: An Example of MQuAKE dataset
Property Value
Edit Request 1 {Lou Pearlman } is a citizen of United States of America $\to$ India
Edit Request 2 The capital of {India} is New Delhi $\to$ Taloga
New Question What is the capital of the country to which Lou Pearlman belonged?
Original Relation (Lou Pearlman, a citizen of, United States of America), (United States of America, the capital of, Washington)
Original Answer Washington
New Relation (Lou Pearlman, a citizen of, India), (India, the capital of, Taloga)
New Answer Taloga
Appendix E Baselines
Our experiments are conducted on GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021) and LLaMA3 (8B) (Dubey et al., 2024). We compare the DiKE against the following state-of-the-art editing techniques:
  • Constrained Fine-Tuning (FT) (Zhu et al., 2020), which directly fine-tunes specific layers of LLM’s parameters using gradient descent while applying regularization constraints to prevent catastrophic forgetting;
  • MEND (Mitchell et al., 2022), a gradient-based low-rank decomposition method that employs a hypernetwork to perform edits;
  • ROME (Meng et al., 2022), which assumes that knowledge in LLMs is stored in FFN modules and performs edits by optimizing and updating specific FFN layers to insert knowledge;
  • MEMIT (Meng et al., 2023), an extension of ROME designed specifically for batch editing tasks by editing a sequence of FFN layers.
  • AlphaEdit (Fang et al., 2025), a null-space projection method designed to better preserve unrelated knowledge during editing by constraining updates orthogonal to preserved information.